# REVEAL-IT: REinforcement learning with Visibility of Evolving Agent poLicy for InTerpretability

### v5

@@ 5.3 Understanding the Relationship between Sub-tasks from Policy
This section will explore how the agent acquires knowledge in various sub-tasks from the perspective of policy. Given that the the different nodes in the policy will be activated in the evaluation, how are the weights linked to these nodes modified? Can GNN-explainer learn to differentiate between the sections that have more significant updates in terms of weight? Fig. 2 illustrates updating the policy during training. Firstly, it is imperative to focus on the modifications in policy updates in every individual row. As training advances, we will observe that the sections with more significant policy updates will undergo modifications (shown by the thick lines). Furthermore, our findings indicate that as the training progresses toward the latter stage, there is a greater overlap between the region with a larger update amplitude and the region triggered by evaluation. This illustrates that GNN explainer comprehends the connection between policy training and evaluation. By carefully considering each vertical direction, we have identified and emphasized the sections of policies that are common to multiple activities (shown by gray boxes). By analyzing the neural network nodes within the gray box, we discovered that certain tasks, which necessitate agents to possess comparable abilities based on common knowledge, exhibit increasingly evident overlaps in their policies. For instance, the subtasks “open microwave 1” and “take apple 1 from microwave 1” both require the agent to comprehend the spatial location of the microwave. However, the second subtask also demands the agent identify the object “apple” and its specific location. There is less collaboration across distinct subtasks. Upon reviewing the prior distance, we observe that the subtask “put apple 1 on sinkbasin 1” shares some similarities with the previous subtask. Both subtasks are interconnected with “apple 1”.
@@ 6 Conclusion and Future Work
We present REVEAL-IT, a novel framework designed to explain the learning process of RL agents in complex environments and tasks while also serving as a foundation for optimizing task sequences. We demonstrate the mechanisms and functionalities of REVEAL-IT in complicated environments. Nevertheless, REVEAL-IT does possess certain constraints. A primary constraint is its ability to adapt to multi-modal challenges. As an illustration, in the ALFWorld, we exclusively implemented the REVEAL-IT within the visual environment. Furthermore, it is worth investigating if REVEAL-IT can effectively transform the knowledge acquired from policies and training problems into natural language, enabling its broader and more direct application in other reinforcement learning domains.
@@ Appendix A Visualized task examples in ALFWorld
Figure 4: Visualized task examples of ALFWorld Shridhar et al. (2021). This benchmark utilizes various household scenarios created within the Ai2Thor environment. In this environment, all objects can be moved to different locations based on available surfaces and class restrictions. This allows for the generation of a wide range of new tasks by combining different objects and goal positions in a procedural manner.
Appendix B More Details about GNN explainer
B.1 GNN explainer highlight the important update
@@
To extract the explanation subgraph $G_{S}$, we first compute the importance weights on edges. A threshold is used to remove low-weight edges, and identify the explanation subgraph $G_{S}$. The ground truth explanations of all datasets are connected subgraphs. Therefore, we identify the explanation as the connected component containing the explained node in GS. We identify the explanation for the generated graph classification by the maximum connected component of $G_{S}$. For all basic RL algorithms, we perform a search to find the maximum threshold such that the explanation is at least of size $K_{M}$. When multiple edges have tied importance weights, all of them are included in the explanation.
B.2 How does GNN explainer map the skills of the RL agent?
@@
In REVEAL-IT, the central concern revolves around the GNN explainer’s capacity to comprehend the agent’s learning prcess and ascertain whether the agent has achieved the certain abilities in accomplishing the task. The GNN explainer is designed to identify and emphasize important updates in the policy. If the nodes associated with these highlighted updates correspond to the nodes engaged by the agent during the evaluation, then it can be inferred that the explanation is cognizant of the specific sections of the policy that align with the agent’s essential abilities. Hence, we ascertain the ratio of nodes identified by the GNN explainer and nodes stimulated by the agent in successful tasks and report it in Fig. 5.
@@
In this series of comparisons, we incorporated a pre-trained Graph Neural Network (GNN) explanation that utilizes data from RL training data from different sequences of training tasks in the same environment. An intriguing observation is that we discovered that the efficacy of the GNN-explainer, when trained on various tasks, was inferior than that of the GNN-explainer taught using a reinforcement learning agent. This discrepancy may arise from the varying demands that different tasks place on the training process of reinforcement learning (RL) agents. Additionally, pre-trained graph neural networks (GNNs) may struggle to accurately comprehend the ongoing RL training progress due to the incorporation of a priori bias.

### v6

@@ 5.3 Understanding the Relationship between Sub-tasks from Policy
This section will explore how the agent acquires knowledge in various sub-tasks from the perspective of policy. Given that the the different nodes in the policy will be activated in the evaluation, how are the weights linked to these nodes modified? Can GNN-explainer learn to differentiate between the sections that have more significant updates in terms of weight? Fig. 2 illustrates updating the policy during training. Firstly, it is imperative to focus on the modifications in policy updates in every individual row. As training advances, we will observe that the sections with more significant policy updates will undergo modifications (shown by the thick lines). Furthermore, our findings indicate that as the training progresses toward the latter stage, there is a greater overlap between the region with a larger update amplitude and the region triggered by evaluation. This illustrates that GNN explainer comprehends the connection between policy training and evaluation. By carefully considering each vertical direction, we have identified and emphasized the sections of policies that are common to multiple activities (shown by gray boxes). By analyzing the neural network nodes within the gray box, we discovered that certain tasks, which necessitate agents to possess comparable abilities based on common knowledge, exhibit increasingly evident overlaps in their policies. For instance, the subtasks “open microwave 1” and “take apple 1 from microwave 1” both require the agent to comprehend the spatial location of the microwave (highlighted by the orange box). However, the second subtask also demands the agent identify the object “apple” and its specific location. There is less collaboration across distinct subtasks. Upon reviewing the prior distance, we observe that the subtask “put apple 1 on sinkbasin 1” shares some similarities with the previous subtask. Both subtasks are interconnected with “apple 1”.
Figure 3: The distribution of verbs in training tasks. This figure demonstrates how REVEAL-IT optimizes the task sequences. We employ various verbs in tasks to differentiate, as each type necessitates distinct capabilities from the agent. The agent training process reflects the task sequence change from left to right. Analyzing the changes in task distribution shows that the “put” type of task is the most prevalent. As the training advances, the initial focus of training tasks is on teaching the agent to locate and retrieve objects in the environment (“look”, “pick”). Subsequently, the agent is trained on tasks that require it to acquire other skills, e.g., “clean”, “heat”, and “examine”. We provide a larger version in Appendix. D.
@@ 6 Conclusion and Future Work
We present REVEAL-IT, a novel framework designed to explain the learning process of RL agents in complex environments and tasks while also serving as a foundation for optimizing task sequences. We demonstrate the mechanisms and functionalities of REVEAL-IT in complex environments. Nevertheless, REVEAL-IT does possess certain constraints. A primary constraint is its ability to adapt to multi-modal challenges. As an illustration, in the ALFWorld, we exclusively implemented the REVEAL-IT within the visual environment. Furthermore, it is worth investigating if REVEAL-IT can effectively transform the knowledge acquired from policies and training problems into natural language, enabling its broader and more direct application in other reinforcement learning domains.
@@ Appendix A Visualized task examples in ALFWorld
Figure 4: Visualized task examples of ALFWorld Shridhar et al. (2021). This benchmark utilizes various household scenarios created within the Ai2Thor environment. All objects can be moved to different locations based on available surfaces and class restrictions in this environment. This allows for the generation of a wide range of new tasks by combining different objects and goal positions in a procedural manner.
Appendix B The difference between REVEAL-IT and other explanation methods
In this section, We want to discuss the novelty of REVEAL-IT with other explanation methods in more detail, and we hope this discussion can help clarify the contribution of REVEAL-IT.
An interactive node-link visualization of convolutional neural networks. Note that utilizing the visualization to show the structure of a CNN/GNN network or an MLP network is previously discussed in  Harley (2015). Nevertheless, the use of visualization is uncommon in RL, this is because:
  • Gaining a visual representation of the policy structure does not enhance our comprehension of the possibilities of RL, and the motivation for visualizing CNNs is essentially distinct from this, i.e., we lack knowledge regarding the specific capabilities of the agent that correspond to a particular section of the policy. REVEAL-IT confirmed the accuracy of this mapping relationship for the first time (refer to Fig. 2).
  • The process of updating policies necessitates millions of revisions. Without a proficient interpreter, meaningful conclusions cannot be derived from the visualization outcomes. The GNN explanation plays a crucial part in REVEAL-IT.
The difference between Gnnexplainer/PGExplainer: Generating explanations for graph neural networks”. The primary objective behind the GNN explainer in REVEAL-IT is to identify the important policy updates that occur during the agent’s learning process. This allows us to correlate these updates with the agent’s abilities, thereby explaining how the agent learns in complex tasks. Simultaneously, the GNN explainer can assist us in determining if the predictor comprehends the inherent connection between the agent’s abilities and the learning process, hence assessing the credibility of the task’s significance. This differs radically from the conventional GNN explainer employed for identifying important edges.
How does REVEAL-IT differ from other interpretable RL algorithms? It is important to understand that REVEAL-IT stands apart from conventional interpretable RL algorithms in: (1) REVEAL-IT does not necessitate the use of artificially constructed causal structures. (2) REVEAL-IT does not require additional knowledge of SCM. These two points are the most significant limitations of classical explanation RL in complex and real environments. To overcome this flaw and restriction, we employ a ”task-agent policy-GNN explainer” architecture, enabling RL use in complex environments and task scenarios. This architecture also comprehensively explains how RL agents learn in these tasks, explaining why an agent can succeed or fail after training.
Appendix C More Details about GNN explainer
C.1 The loss function for GNN explainer
In REVEAL-IT, We use the loss in the PGEexplainer (Luo et al., 2020) to train the GNN explainer, and we use an MSE loss for the GNN predicter. On the other hand, in the traditional GNN explainer method, the learning objectives of the evaluator GNN and the explainer are often consistent, i.e., maximizing the mutual information between the label and the selected subgraph, and we can express this process as:
$\max_{G_{s}}=MI(Y,G_{S})=H(Y)-H(Y|G=G_{S}),$ (3)
where $G_{S}$ denotes the masked subgraph, the GNN explainer optimizes this function locally by training an NN to provide a customized explanation for each instance. However, REVEAL-IT optimizes this by adopting an NN to learn the crucial edges on multiple instances (each evaluation results of RL).
C.2 GNN explainer highlight the important update
@@
To extract the explanation subgraph $G_{S}$, we first compute the importance weights on edges. A threshold is used to remove low-weight edges and identify the explanation subgraph $G_{S}$. The ground truth explanations of all datasets are connected subgraphs. Therefore, we identify the explanation as the connected component containing the explained node in GS. We identify the explanation for the generated graph classification by the maximum connected component of $G_{S}$. For all basic RL algorithms, we perform a search to find the maximum threshold such that the explanation is at least of size $K_{M}$. When multiple edges have tied importance weights, all of them are included in the explanation.
C.3 How to ensure the training process of GNN+RL is stable?
The fundamental principle of the GNN+RL training scheme is mutual learning, where the GNN predicts the value of the training task for the RL agent based on the RL training process, and the GNN predictor influences the training task of the RL agent. During the initial stages of training, the GNN predictor exhibits instability due to inadequate data. We employ a random sampling technique to select training tasks from the complete set. Once RL has accumulated a specific quantity of data, we systematically enhance the significance of the GNN predictor in choosing training tasks. Providing a precise theoretical guarantee for this procedure is challenging, yet it is a commonly used approach in curriculum RL to choose training tasks, e.g.,  Ao et al. (2023); Held et al. (2017).
C.4 How does GNN explainer map the skills of the RL agent?
@@
In REVEAL-IT, the central concern revolves around the GNN explainer’s capacity to comprehend the agent’s learning prcess and ascertain whether the agent has achieved the certain abilities in accomplishing the task. The GNN explainer is designed to identify and emphasize important updates in the policy. Suppose the nodes associated with these highlighted updates correspond to the nodes engaged by the agent during the evaluation. In that case, it can be inferred that the explanation is cognizant of the specific sections of the policy that align with the agent’s essential abilities. Hence, we ascertain the ratio of nodes identified by the GNN explainer and nodes stimulated by the agent in successful tasks and report it in Fig. 5.
@@
In this series of comparisons, we incorporated a pre-trained Graph Neural Network (GNN) explanation that utilizes data from RL training data from different sequences of training tasks in the same environment. An intriguing observation is that we discovered that the efficacy of the GNN-explainer, when trained on various tasks, was inferior than that of the GNN-explainer taught using a reinforcement learning agent. This discrepancy may arise from the varying demands that different tasks place on the training process of reinforcement learning (RL) agents. Additionally, pre-trained graph neural networks (GNNs) may struggle to comprehend the ongoing RL training progress accurately due to the incorporation of a priori bias.
Appendix D Large version of figures in the main paper
Due to the limited space of the main paper, the figures in the main paper might not be clear enough to read. We provide a lager version here for reading.
Figure 6: Visualized important policy updates by GNN explainer. In this figure, the first line depicts the sequential arrangement of sub-tasks that must be accomplished step-by-step to accomplish a given task fully. One of the sub-tasks is located on the left side of the second line. The first to third columns of the tree diagram illustrate the RL policy update process. The blue circles represent nodes in the neural network, while the connections between the circles represent weight updates. Thicker connections indicate larger updates in weight amplitude (selected by GNN explainer). The red circles in the tree diagram on the far right illustrate the active policy nodes during the evaluation process. The links here represent the revised weights throughout the training phase of this subtask. The orange squares indicate the portions of the policy that are common to several sub-tasks. We opt to depict the 8 interconnected nodes with the most significant weight adjustment, facilitating comprehension of the reinforcement learning policy’s learning process and highlighting the policy’s shared component more distinctly.